OpenStudy (anonymous):

find and simplify d/dt [ t^3 e^-2t ].

OpenStudy (anonymous):

answer [t^2 (-2t+3)] / [e^2t]

OpenStudy (cwrw238):

using the product rule; derivative = t^3 * -2 e^-2t + 3t^2 *e^-2t = t^2( -2t + 3) ----------- e^2t yup

OpenStudy (cwrw238):

agreed